OpenStudy (anonymous):

Tyree sold t-shirts for a Math Club fundraiser. During the first week, he sold 9 t-shirts, which was thirty percent of his inventory. How many t-shirts were in his inventory?

OpenStudy (vlery):

30 shirts

OpenStudy (natsent06):

30

OpenStudy (natsent06):

30% = 30 / 100 = 0.3 9 is 30% of what = 9 / 30 = 9 / 0.3 = 30

OpenStudy (natsent06):

did that help you : )

OpenStudy (anonymous):

yes